### v3

# A Forward Reachability Perspective on Control Barrier Functions and Discount Factors in Reachability Analysis

Abstract
Control invariant sets are crucial for various methods that aim to design safe control policies for systems whose state constraints must be satisfied over an indefinite time horizon. In this article, we explore the connections among reachability, control invariance, and Control Barrier Functions (CBFs). Unlike prior formulations based on backward reachability concepts, we establish a strong link between these three concepts by examining the inevitable Forward Reachable Tube (FRT), which is the set of states such that every trajectory reaching the FRT must have passed through a given initial set of states. First, our findings show that the inevitable FRT is precisely this initial set itself if it is a robust control invariant set with a differentiable boundary—a property necessary to connect with CBFs whose zero-level sets are control invariant. We highlight that if the boundary is not differentiable, the FRT of the robust control invariant set may become a strict superset of the invariant set and lose invariance. Next, we formulate a differential game between the control and disturbance, where the inevitable FRT is characterized by the zero-superlevel set of the value function. By incorporating a discount factor in the cost function of the game, the barrier constraint of the CBF naturally arises in the Hamilton-Jacobi equation and determines the optimal policy. Combining these results, the value function of our FRT formulation serves as a CBF-like function, and conversely, any valid CBF is also a forward reachability value function inside the control invariant set, thereby revealing the inverse optimality of the CBF. This strong link between reachability and barrier constraints is not achievable by previous backward reachability-based formulations, and addresses an important gap in existing literature for constructing valid CBFs to ensure safety.
keywords:
Reachability Analysis, Control Invariance, Control Barrier Functions
[…]
spaced
thanks: This work is supported by DARPA Assured Autonomy Grant No. FA8750-18-C-0101, the ONR BRC for Multibody Control Systems No. N000141812214, the NASA ULI Grant No. 80NSSC20M0163, and the NSF Grant CMMI-1931853. Any opinions, findings, and conclusions expressed in this material are those of the authors and do not necessarily reflect the views of any aforementioned organizations.
    * The first two authors contributed equally to this work.
Table 1: Comparison of reachability methods discussed in the paper. More details are discussed in Section 5.
Methods Value function Matching barrier constraint Boundedness Continuity Sol. Unique. of HJ-PDE
[…]
BRT w/o discount [17]
$V(x):=\underset{{\xi_{d}}}{\inf}\;\underset{\mathrm{u}}{\sup}\underset{t\in[0,\infty)}{\inf}h_{S}(\mathrm{x}(t))$
no yes no no
[…]
$\underset{u\in U}{\max}\underset{d\in D}{\min}\frac{\partial V}{\partial x}\cdot f(x,u,d)\geq 0$
BRT with discount [38, 2]
$V(x):=\underset{\xi_{d}}{\inf}\;\underset{\mathrm{u}}{\sup}\underset{t\in[0,\infty)}{\inf}e^{-\gamma t}h_{S}(\mathrm{x}(t))$
no yes yes yes
[…]
$\underset{u\in U}{\max}\underset{d\in D}{\min}\frac{\partial V}{\partial x}\!\cdot\!f(x,u,d)\!-\!\gamma V\!\geq\!0$
CBVF [13, 35]
$V(x):=\underset{\xi_{d}}{\inf}\;\underset{\mathrm{u}}{\sup}\underset{t\in[0,\infty)}{\inf}e^{\gamma t}h_{S}(\mathrm{x}(t))$
yes no no no
[…]
$\underset{u\in U}{\max}\underset{d\in D}{\min}\frac{\partial V}{\partial x}\!\cdot\!f(x,u,d)\!+\!\gamma V\!\geq\!0$
FRT with discount (Ours)
$V(x):=\underset{\xi_{d}}{\sup}\;\underset{\mathrm{u}}{\inf}\underset{t\in(-\infty,0]}{\sup}e^{\gamma t}h_{S}(\mathrm{x}(t))$
yes yes yes yes
[…]
$\underset{u\in U}{\max}\underset{d\in D}{\min}\frac{\partial V}{\partial x}\!\cdot\!f(x,u,d)\!+\!\gamma V\!\geq\!0$
1 Introduction
Safety guarantees are essential for control design in many applications. In this article, we focus on safety problems that can be represented by ensuring that system states satisfy specific constraints over an indefinite time horizon. An effective strategy for ensuring state trajectories stay within the desired constraint region involves identifying a subset where the trajectory can remain for an infinite duration. Sets exhibiting these properties are referred to as control invariant sets [10], and are key to various methods for designing safe control policies [36]. The theoretical analysis of control invariance offers valuable insights for the development of safe control policies.
A typical way of characterizing a control invariant set is by using a scalar function whose zero-superlevel set defines the invariant set, known as the barrier certificate [34]. This concept has evolved into the notion of a control barrier function (CBF) [4], which mandates that the function satisfy a particular differential inequality condition. By enforcing this condition, control policies ensure not just the safety at the boundary of the set, but also a smooth deceleration of trajectories as they approach the boundary. We refer to this condition as the barrier constraint.
This paper delves into control invariance and the barrier constraint using a reachability approach, forging a strong link between reachability, control invariance, and CBFs. We will study systems with disturbance and robust control invariance, control invariance under the worst-case effect of the disturbance as described in [12, 38], and similarly the robust barrier constraint. Reachability analysis forms the basis for many methods that construct robust control invariant sets using dynamic programming principles [9, 28, 17, 2, 38, 13]. These methods compute the invariant sets by eliminating states that will inevitably reach unsafe regions, corresponding to backward reachability problems in which one is concerned with finding states that reach a set of terminal states. However, as we will present, these backward reachability methods only have weak linkage to the CBFs and the barrier constraint.
The key idea of our new formulation is to use the forward reachability concept, in which one is concerned with finding the states that a set of initial states reaches. Our core contribution is a discovery that the optimal policy of the forward reachability problem is determined by the barrier constraint, and that CBFs are forward reachability value functions.
To see this, first, we extend the notion of the minimal forward reachable tube (FRT) [27] to inevitable FRT, which encompasses states that are inevitably reached from the initial set, despite the worst-case disturbance. In this paper, we will use “FRT” to represent this inevitable FRT for brevity, unless specified otherwise. We then verify conditions under which the FRT remains identical to the initial set. We determine that the FRT remains unchanged when the initial set is 1) robustly control invariant, and 2) has a differentiable boundary. It is notable that if either these two conditions is not met, the FRT may differ from the initial set and not be robustly control invariant, whose examples are provided in this article. This result lays the foundation for interpreting CBFs as reachability value functions.
Towards this, we introduce a differential game that characterizes the inevitable FRT as the zero-superlevel set of the value function capturing the game between the control and disturbance. The value function is then determined as a viscosity solution [8] to a particular Hamilton-Jacobi partial differential equation (HJ-PDE) called the Hamilton-Jacobi forward reachable tube variational inequality (HJ-FRT-VI). The crux of this formulation is the incorporation of a discount factor in the cost function of the value function. As a result, first, the barrier constraint emerges in the HJ-FRT-VI, thus becoming the constraint that defines the optimal control policy of the FRT value function. Consequently, the value function acts as a CBF-like function in that it satisfies the barrier constraint almost everywhere. Conversely, any valid CBF also qualifies as a viscosity solution to the HJ-FRT-VI within the control invariant set and can therefore be interpreted as a forward reachability value function.
Additionally, the discount factor induces a contraction mapping in the Bellman operator of the value function, allowing the value function to be continuous and characterized as a unique viscosity solution to the HJ-FRT-VI. While the use of a discount factor in infinite-horizon optimal control problems is not novel [8], we emphasize that our key contribution lies in demonstrating how the discount factor leads to the emergence of the barrier constraint in the reachability formulation.
We highlight that prior formulations relying on backward reachability, introduced in [17, 2, 38, 13], were unable to establish such a connection. For instance, our earlier work in [13] incorporates the barrier constraint into a finite-horizon backward reachability problem, but its extension to the infinite-horizon setting fails to secure the continuity and boundedness of the value function. In contrast, our new formulation is the only method that not only secures the linkage to the barrier constraint but also ensures the boundedness and continuity of the value function and the solution uniqueness of the corresponding HJ-PDE (Table 1). Consequently, by adopting a forward reachability approach to control invariant sets in a manner not previously explored in the literature, our work creates a strong link between these three concepts and thus enables the use of Hamilton-Jacobi methods to characterize CBFs.
The rest of the article is organized as follows. In Section 2, we review the concepts of control invariance and CBFs, extending these notions to systems with disturbances. In Section 3, we introduce forward reachable tubes and present their application to robust control invariant sets. In Section 4, we detail the Hamilton-Jacobi formulation of the FRTs and establish a connection to CBFs. Section 5 compares our formulation and reachability formulations from prior work that have been applied to characterize control invariant sets. We conclude the article with closing remarks in Section 6.
[…]
2 Control Invariance and CBFs
2.1 Control Invariance
We first consider a general nonlinear time-invariant system represented by an ODE
$\dot{\mathrm{x}}(t)=f(\mathrm{x}(t),\mathrm{u}(t))\;\text{for}\;t>0,\qquad\mathrm{x}(0)=x,$ (1)
where $x\in\mathbb{R}^{n}$ is an initial state, $\mathrm{x}:[0,\infty)\rightarrow\mathbb{R}^{n}$ is the solution to the ODE, and $\mathrm{u}:[0,\infty)\rightarrow U$ is a Lebesgue measurable control signal with $U\subset\mathbb{R}^{m_{u}}$. We use $\mathcal{U}$ to denote the set of Lebesgue measurable control signals. We assume that the control input set $U$ is compact, which holds for most physical systems whose actuation limit is bounded. Also, we assume that the system (1) satisfies the following conditions.
Assumption 1 (on vector field of (1)).
  1. 1.
    $f:\mathbb{R}^{n}\times U\rightarrow\mathbb{R}^{n}~{}\textnormal{is uniformly continuous,}$,
  2. 2.
    $f(\cdot,u)$ is Lipschitz continuous in $x\in\mathbb{R}^{n}$ for each $u\in U$,
  3. 3.
    $\exists M>0$ such that $\left\lVert f(x,u)\right\rVert\leq M$ $\forall x\in\mathbb{R}^{n},u\in U$.
Under the above conditions, the solution to the ODE dynamics (1) is unique for any $\mathrm{u}\in\mathcal{U}$ and initial state $x\in\mathbb{R}^{n}$. We will call the solution $\mathrm{x}$ the (forward) trajectory from the initial state $x$.
Let $X\subset\mathbb{R}^{n}$ be the constraint set, i.e. the set that the system must remain within to maintain safety. The main challenge of finding a control signal $\mathrm{u}\in\mathcal{U}$ such that for given $\mathrm{x}(0)\in X$, $\mathrm{x}(t)\in X$ for all $t\geq 0$ (i.e. $\mathrm{x}(\cdot)$ remains safe) is that there may be some states in $X$ from which exiting the set $X$ is inevitable regardless of the choice of $\mathrm{u}$. An effective way of ruling out these failure states is to consider a subset of $X$ that is control invariant.
Definition 1 ((Forward) control invariant [10]).
A set $S\subset\mathbb{R}^{n}$ in the state space is (forward) control invariant under the dynamics (1) if for all $x\in S$, there exists a control signal $\mathrm{u}\in\mathcal{U}$ such that $\mathrm{x}(t)\in S$ for all $t\geq 0$. We also say that such $\mathrm{u}$ renders the trajectory $\mathrm{x}$ forward invariant in $S$.
By the above definition, a trajectory starting inside a control invariant set $S$ that is a subset of $X$ can remain within $S$ for all time, and therefore can stay safe in $X$. The control invariance of a set can be determined by a geometric relationship between the vector field and the tangent cone of the set, as defined next:
Definition 2 ((Bouligand’s) tangent cone [14]).
Given a closed set $S\subset\mathbb{R}^{n}$, the tangent cone to $S$ at $x\in S$ is defined as
$T_{S}(x):=\left\{z\in\mathbb{R}^{n}\mid\liminf_{\tau\rightarrow 0}\frac{\operatorname{dist}(x+\tau z,S)}{\tau}=0\right\},$ (2)
where $\operatorname{dist}(y,S):=\min_{z\in S}{\left\lVert y-z\right\rVert}$.
The tangent cone captures the feasible directions in which one can move from the point $x$ within the set $S$.
Lemma 1.
(Tangential characterization of closed control invariant sets [5, Theorem 11.3.4]) Let the dynamics (1) satisfy Assumption 1. Then, a closed set $S\subset\mathbb{R}^{n}$ is (forward) control invariant under the dynamics (1) if and only if for all $x\in\partial S$,
$\exists u\in U~{}\textnormal{such that}~{}f(x,u)\in T_{S}(x).$ (3)
Below is a corollary of the lemma in a special case when the set $S$ has a differentiable boundary (Assumption 2), by introducing a scalar function $h_{S}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ that satisfies Assumption 3.
Assumption 2.
$S$ is a closed set whose interior is non-empty, and whose boundary, $\partial S$, is continuously differentiable. [^1] For each point $x\in\partial S$, there exists $r>0$ and a $C^{1}$ function $\eta:\mathbb{R}^{n-1}\rightarrow\mathbb{R}$ such that $S\cap B_{r}(x)=\{x\in B_{r}(x)~{}|~{}x_{n}\geq\eta(x_{1},...,x_{n-1})\}$, where relabeling and reorienting the coordinates axes are allowed [15].
Assumption 3.
Given a closed set $S$, $h_{S}:\mathbb{R}^{n}\to\mathbb{R}$ is a function whose zero-superlevel set is $S$, i.e. $S=\{x\in\mathbb{R}^{n}~{}|~{}h_{S}(x)\geq 0\}$, and satisfies the following conditions:
  1. 1.
    $\mathrm{Int}(S)$ =$=\{x\in\mathbb{R}^{n}~{}|~{}h_{S}(x)>0\},$
    $\partial S$ =$=\{x\in\mathbb{R}^{n}~{}|~{}h_{S}(x)=0\}.\vspace{-1em}$ (4)
  2. 2.
    (Differentiability and boundedness) $h_{S}$ is uniformly continuously differentiable and both upper and lower bounded.
  3. 3.
    (Regularity) $\exists{\varepsilon}>0$ such that
    $\frac{\partial h_{S}}{\partial x}(x)\neq 0\quad\forall x\in\partial S+B_{\varepsilon}.\vspace{-0.5em}$ (5)
Lemma 2.
Under Assumptions 1, 2, and 3, $S$ is (forward) control invariant under the dynamics (1) if and only if for all $x\in\partial S$,
$\exists u\in U~{}\textnormal{such that}~{}\frac{\partial h_{S}}{\partial x}(x)\cdot f(x,u)\geq 0.\vspace{-0.5em}$ (6)
Proof.
This is a corollary of Lemma 1 by noticing that for $x\in\partial S$,
$T_{S}(x)=\left\{z\in\mathbb{R}^{n}|\frac{\partial h_{S}}{\partial x}\cdot z\geq 0\right\},\vspace{-0.5em}$ (7)
when Assumptions 2 and 3 hold. ∎
Lemma 2 is known as Nagumo’s theorem for autonomous systems [31]. For a given $S$ that is control invariant, (6) holds for any $h_{S}$ satisfying Assumption 3. The specific choice of $h_{S}$ does not affect condition (6). Also, an $h_{S}$ satisfying Assumption 3 always exists for the set $S$ satisfying Assumption 2, by selecting a regularized distance function for $S$ [25, Theorem 2.1].
Note that a control invariant set does not necessarily have a differentiable boundary. In general, the maximal control invariant set contained in the desired safety constraint set $X$ might have a non-differentiable boundary [5]. However, the differentiability of the boundary will render a few noticeable differences in the theory that will be developed in Section 3, necessitating Assumption 2 for our main theorems.
Next, we consider the concept of control invariance extended to systems with disturbance. There exist various formulations of robustness with respect to disturbances or uncertainties in system dynamics [10, 21, 22, 24]. In this paper, we employ the differential game-based formulation that interprets the disturbance as an adversarial agent playing against the control input [16], as commonly done in the Hamilton-Jacobi analysis for systems with bounded disturbance [12, 17, 38].
For this, we consider the system dynamics
$\dot{\mathrm{x}}(t)=f(\mathrm{x}(t),\mathrm{u}(t),\mathrm{d}(t))\;\text{for}\;t>0,\qquad\mathrm{x}(0)=x,\vspace{-0.5em}$ (8)
where $\mathrm{d}:[0,\infty)\rightarrow D$ is a Lebesgue measurable disturbance signal and $D\subset\mathbb{R}^{m_{d}}$ is a compact set. Note that control systems (1) can be regarded as a special case when the disturbance set $D$ in (8) is set to a singleton (e.g. $D=\{0\}$). We use $\mathcal{D}$ to denote the set of Lebesgue measurable disturbance signals. We assume conditions on the dynamics, similar to Assumption 1:
Assumption 4 (on vector field of (8)).
  1. 1.
    $f:\mathbb{R}^{n}\times U\times D\rightarrow\mathbb{R}^{n}~{}\textnormal{is uniformly continuous,}$,
  2. 2.
    $f(\cdot,u,d)$ is Lipschitz continuous in $x\in\mathbb{R}^{n}$ for each $(u,d)\in U\times D$,
  3. 3.
    $\exists M>0$ such that $\left\lVert f(x,u,d)\right\rVert\leq M$ $\forall x\in\mathbb{R}^{n},(u,d)\in U\times D$,
so that under the above conditions, the solution to the ODE (8) is unique for any pair of $(\mathrm{u},\mathrm{d})\in\mathcal{U}\times\mathcal{D}$ and initial state $x\in\mathbb{R}^{n}$ [16].
To ensure safety under the most adversarial disturbance, we assume that the disturbance can use the control signal’s current and previous information, whereas the control is not aware of the current disturbance input, by considering the notion of the non-anticipative strategies [16]:
$\begin{split}\Xi_{d}\coloneqq\{&\xi_{d}:\mathcal{U}\rightarrow\mathcal{D}~{}|~{}\forall s\in[0,\infty)\text{ and }\mathrm{u},\bar{\mathrm{u}}\in\mathcal{U},\\ &\text{if }\mathrm{u}(\tau)=\bar{\mathrm{u}}(\tau)\text{ a.e. }\tau\in[0,s],\;\\ &\text{then }\xi_{d}[\mathrm{u}](\tau)=\xi_{d}[\bar{\mathrm{u}}](\tau)\text{ a.e. }\tau\in[0,s]\}.\end{split}\vspace{-0.5em}$ (9)
Using the notion of non-anticipative strategies, we define the robust control invariant set under the dynamics (8).
Definition 3 (Robustly (forward) control invariant [12, 38]).
A set $S\subset\mathbb{R}^{n}$ is robustly (forward) control invariant (under the dynamics (8)) if, for all $x\in S$, $\xi_{d}\in\Xi_{d}$, for any $\varepsilon>0$ and $T>0$, there exists a control signal $\mathrm{u}(\cdot)\in\mathcal{U}$ such that $\mathrm{x}(t)\in S+B_{\varepsilon}$ for all $t\in[0,T]$.
If $S$ is an open set, the notion of $\varepsilon$ and $T$ can be dropped. However, at the boundary of a closed set $S$, the disturbance can react to the current control input to drive the system outside of $S$. Thus, $\mathrm{x}$ might not stay in $S$ for all time although the trajectory $\mathrm{x}$ will stay in $S+B_{\varepsilon}$ for any small $\varepsilon$. An example that elucidates the necessity of $\varepsilon$ and $T$ is provided in [12].
Similar to Lemma 2, robustly control invariant sets can be verified by examining the vector field of the dynamics at the boundary of the sets:
Lemma 3.
(Tangential characterization of robustly control invariant sets) Under Assumptions 2, 3, and 4, $S$ is robustly (forward) control invariant under the dynamics (8) if and only if for all $x\in\partial S$,
$\exists u\in U~{}\textnormal{such that}~{}\frac{\partial h_{S}}{\partial x}(x)\cdot f(x,u,d)\geq 0\;\forall d\in D.$ (10)
Proof.
This results from [12, Theorem 2.3], and (7). ∎
Figure 1: (a) For an autonomous system $\dot{\mathrm{x}}(t)=f(\mathrm{x}(t))$, the forward reachable tube (FRT) of the initial set $C$ is the union of the forward trajectories starting from $C$. (b) For control systems, the viable (maximal) FRT is the collection of all possible trajectories that depart from $C$. (c) On the other hand, the inevitable (minimal) FRT, the main focus of this study, is a collection of states such that every trajectory reaching it must have passed through $C$ at some point in the past.
2.2 Control Barrier Functions
Lemma 2 implies that a safe control input on the boundary of the set $S$ can render the trajectory $\mathrm{x}$ forward invariant in $S$. Control barrier functions (CBFs), first introduced in [4], additionally impose conditions on the input when the trajectory is strictly inside the set before reaching the boundary, which enables the trajectories to “smoothly brake” as they approach the boundary. Here, we extend its definition to robust CBF for the dynamics with disturbance (8).
Definition 4 (Robust Control Barrier Function).
A function $h_{S}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ that satisfies Assumption 3 for a closed set $S$ is a robust CBF for the dynamics (8) if there exists an extended class $\mathcal{K}$ function $\alpha$ such that, for all $x\in S$,
$\max_{u\in U}\min_{d\in D}\frac{\partial h_{S}}{\partial x}(x)\cdot f(x,u,d)+\alpha\left(h_{S}(x)\right)\geq 0.$ (11)
Here, $\alpha:\mathbb{R}\rightarrow\mathbb{R}$ is an extended class $\mathcal{K}$ function if it is continuous and strictly increasing and satisfies $\alpha(0)=0$. We say the barrier constraint is feasible at $x$ if condition (11) holds for $x$.
This paper considers a particular class $\mathcal{K}$ function $\alpha(y)=\gamma y$ for a constant $\gamma>0$, as in [33, 32],
$\max_{u\in U}\min_{d\in D}\frac{\partial h_{S}}{\partial x}(x)\cdot f(x,u)+\gamma h_{S}(x)\geq 0.$ (12)
This is the most common choice of class $\mathcal{K}$ function used in the CBF literature, and enables us to make a connection between CBFs and reachability value functions where $\gamma$ will play the role of a discount factor in the reachability formulation. Intuitively, (12) ensures that $h_{S}(\mathrm{x}(t))$ does not decay faster than the exponentially decaying curve $\dot{h}_{S}(\mathrm{x}(t))=-\gamma h_{S}(\mathrm{x}(t))$. This induces the braking mechanism to any trajectory $\mathrm{x}$ approaching the boundary of $S$. If (12) is satisfied, (10) is trivially satisfied at $x\in\partial S$ where $h_{S}(x)=0$. Thus, according to Lemma 3, the existence of the robust CBF $h_{S}$ is a sufficient condition for $S$ being robust control invariant.
Proposition 1.
Let the dynamics (8) satisfy Assumption 4 and let $S\subset\mathbb{R}^{n}$ satisfy Assumption 2. If a robust CBF $h_{S}$ exists, $S$ is robustly control invariant.
Remark 1.
For systems without disturbance (1), if (6) in Lemma 2 holds with strict inequality for a compact and control invariant set $S$, any $h_{S}$ satisfying Assumption 3 is a CBF with large enough $\gamma$ [4, Lemma 2].
3 Forward Reachability Analysis
In this section, we apply forward reachability analysis to control invariant sets. In order to do so, we first provide background on the forward reachability of a set and introduce definitions of forward reachable tubes.
3.1 Forward Reachability
Forward reachability analyzes a set’s evolution in the future; its purpose is to identify the states that trajectories from an initial set $C\subset\mathbb{R}^{n}$ reach forward in time. The forward reachable tube (FRT) of a set, roughly speaking, encompasses states that are reached by trajectories that depart from the initial set. This concept is illustrated in Figure 1. For autonomous systems (e.g. $\dot{\mathrm{x}}(t)=f(\mathrm{x}(t))$), the forward evolution of a set is uniquely determined (Fig 1 (a)). However, for systems with control and/or disturbance inputs like (1) and (8), the trajectory, and thus the forward reachable tube, can be determined in various ways according to control and disturbance.
Consider the dynamics without disturbance, satisfying (1). At one extreme, the control can use its best effort to get further away from the original set $C$, and at the other extreme, the control works to stay as close to $C$ as possible. The former would cause the FRT to expand maximally covering all the states such that from $C$, reaching them is viable (Fig 1 (b)), and the latter would induce the FRT to grow minimally, encapsulating only the states which inevitably must have evolved from $C$ (Fig 1 (c)). From this intuition, we are able to define the viable (maximal) and inevitable (minimal) FRTs of a set $C$. Here, we only define the inevitable FRT which is the focus of this study, and readers are referred to [27] for the definition of the viable FRT.
Applications of forward reachability in the safe control and verification literature primarily focused on determining the viable FRT of the set of initial states, and checking whether this set intersects with the unsafe set [3, 23, 37]. In this context, the utility of inevitable FRTs would be limited, since the trajectories from the initial set could still enter the unsafe region even if the inevitable FRT does not [27]. In this paper, we revisit the utility of the inevitable FRT under the specific context in which the initial set is robustly control invariant.
To introduce the formal definition of the inevitable FRT, we use a separate notation for the solution of the ODE whose terminal state is specified as $x$ (as opposed to initial states being specified as $x$ in (8)):
$\dot{\mathrm{x}}^{-}(t)=f(\mathrm{x}^{-}(t),\mathrm{u}^{-}(t),\xi_{d}^{-}[\mathrm{u}^{-}](t)),\;t<0,\;\;\mathrm{x}^{-}(0)=x,$
where $\mathrm{u}^{-}:(-\infty,0]\rightarrow U$ is an element of $\mathcal{U}^{-}$, a set of measurable backward control signal. With $\mathcal{D}^{-}$ denoting a set of measurable backward disturbance signals $\mathrm{d}^{-}:(-\infty,0]\rightarrow D$, $\xi_{d}^{-}$ is a non-anticipative strategy for the disturbance backward in time. The separate notation is necessary due to the causality of the disturbance strategy being reversed in time. We will call $\mathrm{x}^{-}$ the backward trajectory of (terminal state) $x$. Evaluating whether $\mathrm{x}^{-}$ reaches the set $C$ when time is considered to flow backward will tell us whether $x$ belongs to the forward reachable tube of $C$.
Definition 5 (Inevitable FRT).
For a given initial set $C\subset\mathbb{R}^{n}$ which is an open set, we define the (infinite-horizon inevitable) FRT of $C$ as the following set.
$\mathrm{FRT}(C)\coloneqq\Big{\{}x\!\in\!\mathbb{R}^{n}~{}\Big{|}~{}\exists\xi_{d}^{-}\in\Xi_{d}^{-},T>0\text{ s.t.}\forall\mathrm{u}\in\mathcal{U}^{-},\;\;\;\;\;\;$
[…]
Note that  $\mathrm{FRT}(\cdot)$ can be interpreted as a set mapping, $\text{FRT}:2^{\mathbb{R}^{n}}\rightarrow 2^{\mathbb{R}^{n}}$. In words, $\mathrm{FRT}(C)$ is a collection of states such that every trajectory reaching it forward in time must have passed through $C$ at some point in the past. The FRT is shaped by the control aiming to restrain the growth of the FRT, whereas the disturbance is assumed to act adversarially and attempts to grow the FRT. For control systems, the inevitable FRT is minimal [27] since it excludes any state that can be reached by a trajectory that does not evolve from $C$, which is particularly relevant to the concept of control invariance.
3.2 Forward Reachable Tubes of Control Invariant Sets
We now investigate the relationship between the inevitable FRTs and control invariance. The main theorem of the section is as follows:
Theorem 1.
Suppose $f$ satisfies Assumption 4. Then, a set $S$ satisfying Assumption 2 is robustly control invariant under (8) if and only if $\mathrm{FRT}(\mathrm{Int}(S))=\mathrm{Int}(S)$.
The theorem identifies that the interior of any control invariant set with a differentiable boundary is a fixed point of $\mathrm{FRT}(\cdot)$. The non-triviality of the theorem arises from the need for not only control invariance but also the differentiability of the boundary of the set $S$.
Note that, in contrast to the forward reachability result in Theorem 1, backward reachability characterizes closed robust control invariant sets as fixed points of the viability kernel operation [5], without requiring the invariant set to have a differentiable boundary. Nevertheless, our motivation for focusing on forward reachability lies in its potential to reveal connections between reachability value functions and CBFs—a connection that will be explored in the subsequent sections. We will show that backward reachability value functions are fundamentally distinct from CBFs, limiting their utility for safe control synthesis.
The rest of this section presents the proof of Theorem 1. For the proof, we have to reason about the backward flow of the dynamics, as in (13). Thus, we first consider the notion of backward control invariance, the mirrored version of the forward control invariance.
Definition 6 (Robustly backward control invariant).
A set $S\in\mathbb{R}^{n}$ is robustly backward control invariant (under (8)) if for all $x\in S$, for all $\xi_{d}^{-}\in\Xi_{d}^{-}$, for any $\varepsilon>0$ and time $T>0$, there exists a backward control signal $\mathrm{u}^{-}\in\mathcal{U}^{-}$ such that $\mathrm{x}^{-}(t)\in S+B_{\varepsilon}$ for all $t\in[-T,0]$, where $\mathrm{x}^{-}$ solves (13).
Put simply, backward control invariant sets are forward control invariant under the negated dynamics (where the time flows inversely). We can characterize the backward control invariant sets similarly to Lemma 3:
Corollary 1.
Under Assumptions 2, 3, and 4, $S$ is robustly backward control invariant if and only if for all $x\in\partial S$,
$\exists u\in U~{}\textnormal{s.t.}~{}-\frac{\partial h_{S}}{\partial x}(x)\cdot f(x,u,d)\geq 0\;\forall d\in D.\vspace{-1em}$ (15)
By combining Lemma 3 and Corollary 1, we draw a connection between forward and backward invariant sets.
Figure 2: Forward reachable tubes under double integrator dynamics for various shapes of $S$. In the first row, the initial set $S$ is visualized as the interior of the pink level curve. The interior of the blue level curve in the second row is $\mathrm{FRT}(\mathrm{Int}(S))$ for each case. (a) Smooth $S_{a}$ that is not control invariant, resulting in $\mathrm{FRT}(\mathrm{Int}(S_{a}))\neq\mathrm{Int}(S_{a})$. The FRT is still not control invariant. (b, c) Nonsmooth sets $S_{b}$, $S_{c}$ that are control invariant; in case (b), $\mathrm{FRT}(\mathrm{Int}(S_{b}))\neq\mathrm{Int}(S_{b})$, and in case (c), $\mathrm{FRT}(\mathrm{Int}(S_{c}))=\mathrm{Int}(S_{c})$. This shows that Assumption 2 is required for Theorem 1 to hold. (d) Smooth control invariant $S_{d}$ that results in $\mathrm{FRT}(\mathrm{Int}(S_{d}))=\mathrm{Int}(S_{d})$ according to Theorem 1.
Lemma 4.
Under Assumptions 2, 3, and 4, $S$ is robustly forward control invariant if and only if $\mathrm{Int}(S)^{c}$ is robustly backward control invariant.
Proof.
By Lemma 3, $S$ is robustly forward control invariant if and only if for all $x\in\partial S$, (10) is satisfied. Note that $\partial S=\partial\mathrm{Int}(S)^{c}$ and $\mathrm{Int}(S)^{c}$ and $h_{\mathrm{Int}(S)^{c}}:=-h_{S}$ also satisfies Assumptions 2 and 3, respectively. Since $\frac{\partial h_{\mathrm{Int}(S)^{c}}}{\partial x}(x)=-\frac{\partial h_{S}}{\partial x}(x)$, (10) is equivalent to
$\exists u\in U~{}\textnormal{s.t.}~{}\frac{\partial h_{\mathrm{Int}(S)^{c}}}{\partial x}(x)\cdot\left(-f(x,u,d)\right)\geq 0\;\forall d\in D.$
By applying Corollary 1, $\mathrm{Int}(S)^{c}$ is robustly backward control invariant if and only if $\forall x\!\in\!\partial S$, (16) holds. ∎
In Lemma 4, Assumption 2 guarantees that, for a state $x_{1}$ on the boundary of $S$, if there exists a particular control $u_{1}$ such that $f(x_{1},u_{1},d)$ points inward to $S$ for all $d\in D$, $-f(x_{1},u_{1},d)$ points outwards to $S$ for all $d\in D$.
Next, we introduce the concept of a viability kernel under the backward dynamics:
Definition 7.
A viability kernel of a closed set $C\subset\mathbb{R}^{n}$ under the backward dynamics (13), is defined as:
[…]
This definition applies the concept of leadership kernel in [12] to the backward dynamics. Given any closed set $C\subset\mathbb{R}^{n}$, the $\mathrm{VK}^{-}(C)$ is the largest negatively robustly invariant set in $C$. Thus, if $C$ is negatively robustly control invariant, $C$ is identical to $\mathrm{VK}^{-}(C)$:
Lemma 5.
A closed set $C\subset\mathbb{R}^{n}$ is negatively robustly control invariant under (8) if and only if $\mathrm{VK}^{-}(C)=C$.
Proof.
This is a direct results of the definitions. ∎
The FRT and the viability kernel under the backward dynamics have the following complement property.
Lemma 6.
For any open set $C\subset\mathbb{R}^{n}$, the state space $\mathbb{R}^{n}$ can be partitioned into $\mathrm{FRT}(C)$ and $\mathrm{VK}^{-}(C^{c})$, i.e.,
$\{\mathrm{FRT}(C)\}^{c}=\mathrm{VK}^{-}(C^{c}).\vspace{-0.5em}$ (17)
Furthermore, $\mathrm{VK}^{-}(C)$ is always a closed set and $\mathrm{FRT}(C)$ is always an open set.
Proof.
The complement relationship (17) follows directly from Definitions 5 and 7. That $\mathrm{VK}^{-}(C)$ is a closed set is proven in [12], and $\mathrm{FRT}(C)$ being an open set results from the complement relationship. ∎
Based on the lemmas above, we present the proof of Theorem 1 below.
Proof.
(Theorem 1) $S$ is robustly forward control invariant if and only if $\mathrm{Int}(S)^{c}$ is robustly backward control invariant, by Lemma 4. By Lemma 5, $\mathrm{Int}(S)^{c}$ is robustly backward control invariant if and only if $\mathrm{VK}^{-}(\mathrm{Int}(S)^{c})=\mathrm{Int}(S)^{c}$. By Lemma 6, $\mathrm{VK}^{-}(\mathrm{Int}(S)^{c})=\mathrm{FRT}(\mathrm{Int}(S))^{c}$. Thus, from the above statements, $S$ is robustly forward control invariant if and only if $\mathrm{FRT}(\mathrm{Int}(S))=\mathrm{Int}(S)$. ∎
The differentiability of the boundary (Assumption 2) is necessary for the crucial step of the theorem’s proof, which uses the equivalence between $f(x,\bar{\pi}(x))\in T_{S}(x)$ and $-f(x,\bar{\pi}(x))\in T_{\mathrm{Int}(S)^{c}}(x)$ (Lemma 4). At the non-smooth boundary of $S$, this is not necessarily true. In this case, a forward trajectory can be inevitably “leaked” from the interior of $S$, leading to the expansion of $\mathrm{FRT}(\mathrm{Int}(S))$ to a strict superset of $\mathrm{Int}(S)$. An example of this incident is introduced next (Figure 2b).
3.3 Example: Double Integrator
We introduce examples of four different initial sets (Figure 2 first row) and their FRTs (Figure 2 second row) to illustrate the importance of two key assumptions—control invariance of the initial set and differentiability of its boundary—for ensuring that FRTs are themselves control invariant. These examples are based on a simple double integrator system defined by $\dot{p}=v,\dot{v}=u$, with state $x=[p\;v]^{T}$, and control input $u$, with control bound $u\in[-1,1]$. Note that curves $p=c\pm\frac{1}{2}v^{2}$ characterize trajectories that decelerate or accelerate until $v=0$, with the saturated input, $u=\pm 1$. The four specific $S$ sets are:
  1. (a)
    $S_{a}$ is defined as the circular region with radius $r$ centered at the origin. This set $S_{a}$ satisfies Assumption 2 but is not control invariant.
  2. (b)
    $S_{b}$ is formed by five curves, $p=-p_{1}+\frac{1}{2}v^{2}$, $p=p_{1}-\frac{1}{2}v^{2}$, $p=p_{2}-\frac{1}{2}v^{2}$, $p=-p_{2}+\frac{1}{2}v^{2}$, $p=-\frac{1}{2}v^{2}$, and the $p$-axis, as shown in Figure 2(b). This set does not satisfy Assumption 2, but is control invariant.
  3. (c)
    $S_{c}$ is formed by two curves, $p=-p_{1}+\frac{1}{2}v^{2}$, and $p=p_{1}-\frac{1}{2}v^{2}$. This set also does not satisfy Assumption 2 but is control invariant.
  4. (d)
    $S_{d}$ is formed by two curves, $p=-p_{1}+\frac{1}{2}v^{2}$ and $p=p_{1}-\frac{1}{2}v^{2}$ ($p_{1}>1$), and two arcs whose radius is $r=-1+2\sqrt{p_{1}}$ that are tangential to the curves whose centers are positioned at $(-p_{1}+r,0),(p_{1}-r,0)$, respectively. This set satisfies Assumption 2 and is also control invariant.
The first case demonstrates that the FRT can be a strict superset of $\mathrm{Int}(S_{a})$ when $S_{a}$ is not control invariant. The resulting $\mathrm{FRT}(\mathrm{Int}(S_{a}))$ is still not control invariant since the trajectory is bound to escape the set at points A and B. This example reveals a challenge in constructing a control invariant set with forward reachability when the initial set is not control invariant. In the second and third cases, the control invariance of $S$ can be checked analytically. The second case shows that the FRT can be a strict superset of $\mathrm{Int}(S_{b})$ if Assumption 2 is not met, even though $S_{b}$ is control invariant. Note that point C is where $f(x,\bar{\pi}(x))\!\in\!T_{S}(x)$ holds but $-f(x,\bar{\pi}(x))\!\in\!T_{\mathrm{Int}(S)^{c}}(x)$ does not hold. In the third case, since $-f(x,\bar{\pi}(x))\!\in\!T_{\mathrm{Int}(S)^{c}}(x)$ holds at both points D and E where $\partial S_{c}$ is not smooth, $\mathrm{FRT}(\mathrm{Int}(S_{c}))=\mathrm{Int}(S_{c})$. Finally, the set $S_{d}$ in the last case satisfies Assumption 2 and is also control invariant. Thus, according to Theorem 1, $\mathrm{FRT}(\mathrm{Int}(S_{d}))$ remains the same as $\mathrm{Int}(S_{d})$.
4 FRT value function and CBF
In this section, by taking the Hamilton-Jacobi approach to the forward reachability problem, we pose the computation of FRT as a differential game. The FRT is characterized by the value function proposed in Section 4.1. This value function is the unique solution to the HJ-PDE proposed in Section 4.2. In Section 4.3, under the assumption that the initial set is control invariant and has a differentiable boundary, we establish a connection between the proposed value function and the CBFs. Importantly, this provides an interpretation of any valid CBF as a forward reachability value function.
4.1 FRT Value Function with Discount Factor
By noting that $h_{S}$ satisfying Assumption 3 serves as a distance-like metric to the boundary of $S$ and its sign serves as an indicator of the inclusion in $S$, we can rewrite the definition of the FRT in (14) as follows:
$\mathrm{FRT}(\mathrm{Int}(S))$
=$=\!\big{\{}x~{}|~{}\exists\xi_{d}^{-}\in\Xi_{d}^{-},\forall\mathrm{u}^{-}\!\in\!\mathcal{U}^{-},\!\sup_{t\in(-\infty,0]}h_{S}(\mathrm{x}^{-}(t))\!>\!0\big{\}}$
=$=\!\big{\{}x~{}|~{}\sup_{\xi_{d}^{-}\in\Xi_{d}^{-}}\inf_{\mathrm{u}^{-}\in\mathcal{U}^{-}}\sup_{t\in(-\infty,0]}h_{S}(\mathrm{x}^{-}(t))>0\big{\}}.$
Since rescaling $h_{S}(\mathrm{x}^{-}(t))$ with a positive constant at any time $t$ does not change its sign, the following holds:
$\mathrm{FRT}(\mathrm{Int}(S))\!=\Big{\{}x~{}|\sup_{\xi_{d}^{-}\in\Xi_{d}^{-}}\inf_{\mathrm{u}^{-}\in\mathcal{U}^{-}}\sup_{t\in(-\infty,0]}e^{\gamma t}h_{S}(\mathrm{x}^{-}(t))>0\Big{\}},$
where $\gamma\!>\!0$ and at each time $t\!\in\!(-\infty,0]$, $h_{S}(\mathrm{x}^{-}(t))$ is rescaled by $e^{\gamma t}$. Thus, by defining the FRT value function of $S$, $V_{\gamma}:\!\mathbb{R}^{n}\!\rightarrow\!\mathbb{R}$, as
$V_{\gamma}(x)\coloneqq\sup_{\xi_{d}^{-}\in\Xi_{d}^{-}}\inf_{\mathrm{u}^{-}\in\mathcal{U}^{-}}J_{\gamma}(x,\mathrm{u}^{-},\Xi_{d}^{-})\vspace{-1em}$ (18)
with the cost functional $J_{\gamma}\!:\!\mathbb{R}^{n}\!\times\!\mathcal{U}^{-}\!\times\Xi_{d}^{-}\!\rightarrow\!\mathbb{R}$ defined as
$J_{\gamma}(x,\mathrm{u}^{-},\xi_{d}^{-})=\sup_{t\in(-\infty,0]}e^{\gamma t}h_{S}(\mathrm{x}^{-}(t)),\vspace{-1em}$ (19)
where $\mathrm{x}^{-}$ solves (13) and $x$ is the terminal state of $\mathrm{x}^{-}$, the following holds.
Lemma 7.
Suppose $S\in\mathbb{R}^{n}$ is a closed set, $f$ satisfies Assumption 4, and a bounded function $h_{S}$ satisfies Assumption 3-1). $V_{\gamma}(x)$ is positive if and only if $x$ belongs to the FRT of the interior of $S$:
$\mathrm{FRT}(\mathrm{Int}(S))=\{x~{}|~{}V_{\gamma}(x)>0\}.$ (20)
Proof.
See Appendix 7.1. ∎
The value function (18) captures a differential game between the control and the disturbance, wherein the optimal control signal of this game is verifying the existence of a trajectory that reaches $x$ without passing through $\mathrm{Int}(S)$ in the past under the worst-case disturbance. If such a trajectory does not exist, $V_{\gamma}(x)$ is positive and $x$ is inside $\mathrm{FRT}(\mathrm{Int}(S))$.
We now discuss the effect of introducing $\gamma$ to the cost function. (18) defines a differential game with a discounted supremum-over-time cost function. Larger $\gamma$ will recognize the value of $h_{S}$ at the current time more than the value in the past. More importantly, $\gamma$ would also affect the resulting optimal control policy; whereas when there is no discount, the optimal control always has to try its best to maintain the value of $h_{S}$, the discount alleviates this conservativeness and allows the optimal control to decay the value of $h_{S}$.
As such, the discount factor introduces the “game-of-degree” aspect to the reachability problem. In this game-of-degree, the parameter $\gamma$ serves as a knob that adjusts how conservative the resulting optimal policy will be. However, the fundamental nature of the reachability problem, what is called the “game-of-kind” [12, 6]—whether or not a state is inside the FRT—remains consistent, since (20) holds for any value of $\gamma$ and the property can be determined by the positivity of $V_{\gamma}(x)$.
The introduction of a discount in (19) is similar to introducing a discount to infinite-horizon sum-over-time cost optimal control problems [8]. In fact, many favorable properties of the value function resulting from the discount, including its Lipschitz continuity and the contraction of the corresponding Bellman backup operator, hold similarly.
Proposition 2 (Lipschitz Continuity).
Suppose $f$ satisfies Assumption 4 and $h_{S}$ is Lipschitz continuous. $V_{\gamma}$ is Lipschitz continuous in $\mathbb{R}^{n}$ if $L_{f}<\gamma$, where $L_{f}$ is the Lipschitz constant of $f$.
Proof.
See Appendix 7.2. ∎
The condition $L_{f}<\gamma$ implies that the discount factor has to be large enough to suppress the effect of the vector field in prohibiting continuity. Under this condition, since the value function is Lipschitz continuous, it is differentiable almost everywhere by Rademacher’s Theorem. As we will later see, other infinite-horizon value functions in backward reachability formulations [17, 13] do not have Lipschitz continuity and can even be discontinuous, which prohibits the usage of a differential inequality-based condition like the barrier constraint for deriving safe control policies from the value function.
The contraction property of the Bellman backup will be discussed next after introducing the HJ-PDE characterization of $V_{\gamma}$, which provides a computational machinery for the computation of the value function $V_{\gamma}$.
4.2 Hamilton-Jacobi Characterization
The HJ-PDE underlying the FRT value function $V_{\gamma}$ is derived by applying Bellman’s principle of optimality to (18), which results in the following variational inequality.
Theorem 2 (Forward Reachable Tube Hamilton-Jacobi Variational Inequality).
Suppose $h_{S}$ is a bounded and Lipschitz continuous function, and $\gamma>0$. $V_{\gamma}$ in (18) is a unique viscosity solution [8] in $\mathbb{R}^{n}$ of the following HJ-PDE, called forward reachable tube Hamilton-Jacobi variational inequality (FRT-HJ-VI):
[…]
$\max_{u}\min_{d}\frac{\partial V_{\gamma}}{\partial x}\cdot f(x,u,d)+\gamma V_{\gamma}(x)\Big{\}}.\vspace{-0.5em}$
Proof.
See Appendix 7.3. ∎
For non-positive values of $\gamma$, $V_{\gamma}$ might be unbounded and the FRT-HJ-VI might have multiple solutions. In contrast, a strictly positive value of $\gamma$ guarantees the boundedness and the uniqueness of the solution of the FRT-HJ-VI. (An example in Appendix illustrates these outcomes.) In fact, the uniqueness property follows from the contraction property of the Bellman backup associated with the dynamic programming principle of $V_{\gamma}$.
To see this, we define a Bellman backup operator $B_{T}:\text{BUC}(\mathbb{R}^{n})\rightarrow\text{BUC}(\mathbb{R}^{n})$ for $T>0$, where $\text{BUC}(\mathbb{R}^{n})$ represents a set of bounded and uniformly continuous functions: $\mathbb{R}^{n}\rightarrow\mathbb{R}$, as
[…]
$\!\!\!e^{-\gamma T}V(\mathrm{x}^{-}(-T))\Big{\}}.\vspace{-1em}$ (22)
Then, the following holds.
Theorem 3 (Contraction mapping).
For $V^{1},V^{2}\!\in\!\text{BUC}(\mathbb{R}^{n})$,
$\lVert B_{T}[V^{1}]-B_{T}[V^{2}]\rVert_{\infty}\leq e^{-\gamma T}\lVert V^{1}-V^{2}\rVert_{\infty},\vspace{-0.75em}$ (23)
and the FRT value function $V_{\gamma}$ in (18) is the unique fixed-point solution to $V_{\gamma}\!=\!B_{T}[V_{\gamma}]$ for each $T\!>\!0$. Also, for any $V\!\in\!\textnormal{BUC}(\mathbb{R}^{n})$, $\lim_{T\rightarrow\infty}B_{T}[V]=V_{\gamma}$.
Proof.
See Appendix 7.4. ∎
Theorem 3 provides various ways to compute the FRT value function $V_{\gamma}$ using the operation $B_{T}[\cdot]$, which do not require any assumptions for the initial guess of the value function, besides the boundedness and uniform continuity in $\mathbb{R}^{n}$. For instance, it allows the computation of $V_{\gamma}$ with a finite-horizon HJ-PDE, for which numerical solution methods like the level set method [29] are well established. The finite-horizon value function is guaranteed to converge to $V_{\gamma}$ as $T\rightarrow\infty$. Also, the theorem enables other numerical schemes based on time discretization, such as value iteration. For further details of how Theorem 3 can be utilized for finite-horizon-based computation or value iteration, see Appendix 7.5.
4.3 FRT Value Functions for Control Invariant Sets
We now revisit the forward reachability for control invariant sets, extending the analysis in Section 3.2, and draw a connection between the FRT value function $V_{\gamma}$ and robust CBFs.
First, when $S$ has a differentiable boundary and is robustly control invariant where Theorem 1 holds, the following holds:
Proposition 3 (FRT value function characterization of robust control invariant sets).
Under Assumptions 2, 3, and 4, $S$ is robustly control invariant if and only if
$\mathrm{FRT}(\mathrm{Int}(S))=\mathrm{Int}(S)\;$ =$=\{x~{}|~{}V_{\gamma}(x)>0\},$
$\mathrm{Int}(S)^{c}$ =$=\{x~{}|~{}V_{\gamma}(x)=0\},$
where $V_{\gamma}$ is defined as (18).
Proof.
This results from Theorem 1 and Lemma 7. ∎
When Proposition 3 holds, the control invariant set $\mathrm{Int}(S)$ is characterized as a strict zero-superlevel of $V_{\gamma}$. This enables the synthesis of a control policy using $V_{\gamma}$ to maintain forward invariance of trajectories within $\mathrm{Int}(S)$. To see this, we derive an optimal policy of $V_{\gamma}$ from the FRT-HJ-VI (21).
Proposition 4 (Optimal policy of $V_{\gamma}$).
Under the assumptions in Theorem 2, we define the set-valued map policy $K_{\gamma}:S\rightarrow 2^{U}$ as
[…]
where $V_{\gamma}$ is defined as (18). Then, $K_{\gamma}(x)$ is non-empty for every $x\in\mathrm{Int}(S)$ where $\frac{\partial V_{\gamma}}{\partial x}$ exists. In addition, if $V_{\gamma}$ is differentiable, any element of $K_{\gamma}(x)$ is an optimal control input with respect to $V_{\gamma}$ in (18), and under Assumptions 2, 3, and 4, if $S$ is robustly control invariant, the trajectory under $K_{\gamma}(x)$ remains forward invariant in $S$ under the worst-case disturbance.
Proof.
See Appendix 7.6. ∎
Note that the non-emptiness of $K_{\gamma}(x)$ is derived from $V_{\gamma}$ satisfying the FRT-HJ-VI (21). By noting that the second term of the minimum in (21) has to be non-negative for (21) to hold, we get that
$\max_{u\in U}\min_{d\in D}\frac{\partial V_{\gamma}}{\partial x}\cdot f(x,u,d)+\gamma V_{\gamma}(x)\geq 0\vspace{-0.5em}$ (25)
at every $x\in\mathbb{R}^{n}$ where $V_{\gamma}$ is differentiable. This corresponds to the barrier constraint in (11) where $\alpha(y)=\gamma y$.
Since the value function is Lipschitz continuous and differentiable almost everywhere by Proposition 2, $V_{\gamma}$ satisfies (25) almost everywhere in $\mathrm{Int}(S)\subset\mathbb{R}^{n}$. Note that $V_{\gamma}$ is 0 everywhere outside $\mathrm{Int}(S)$. If $V_{\gamma}$ is differentiable in $\mathrm{Int}(S)$, (25) is satisfied everywhere in $\mathrm{Int}(S)$, which constitutes the robust CBF in Definition 4:
Corollary 2.
If $V_{\gamma}$ is continuously differentiable in $\mathrm{Int}(S)$, $V_{\gamma}:S\rightarrow\mathbb{R}$ is a robust CBF.
Proof.
Proposition 4 implies that the barrier constraint holds at any state in $\mathrm{Int}(S)$. As we constrain the domain of $V_{\gamma}$ to $S$, the gradient of $V_{\gamma}$ at $x\in\partial S$, is defined as $\lim_{y\rightarrow x}\frac{\partial V_{\gamma}}{\partial x}$. Since $V_{\gamma}$ is continuously differentiable, this limit exists and $K_{\gamma}(x)$ in (24) is nonempty by Proposition 4. Thus, the statement holds by the definition of robust CBF in Def. 4. ∎
More importantly, any valid robust CBF $h$ itself is the FRT value function in $\mathrm{Int}(S)$:
Theorem 4 (Inverse optimality of CBFs).
Under Assumptions 2 and 4, let $h:\mathbb{R}^{n}\rightarrow\mathbb{R}$ be a differentiable function that satisfies Assumption 3 and is a robust CBF for a closed set $S$, satisfying
$\max_{u\in U}\min_{d\in D}\frac{\partial h}{\partial x}\cdot f(x,u,d)+\gamma h(x)\geq 0,\vspace{-0.75em}$ (26)
for all $x\in S$ and some $\gamma>0$. Then,
$V_{\gamma}(x)=\max\{0,h(x)\}\vspace{-0.75em}$ (27)
is the unique viscosity solution of the FRT-HJ-VI (21) with $h_{S}(x)=h(x)$.
Proof.
See Appendix 7.7. ∎
Corollary 2 and Theorem 4 establish a tight theoretical linkage between HJ reachability analysis and CBFs, wherein the role of discount factor is crucial. By introducing the discount factor to the reachability formulation, the value function becomes a CBF-like function in that it satisfies the barrier constraint almost everywhere in the set $S$, and in the best case when it is differentiable, it becomes the CBF. Moreover, by Proposition 4, the barrier constraint defines the optimal policy. On the other hand, by Theorem 4, any CBF can be interpreted as an FRT value function with a discount factor. Thus, the inverse optimal control principle [1] underlying the barrier constraint and the CBF is the discounted forward reachability. The discounted FRT cost function (19) is the inverse optimal cost that characterizes the CBF itself as the value function, and any control satisfying the barrier constraint as the corresponding optimal control. In [24], the inverse optimality of CBF-based min-norm controllers has been investigated as an infinite-horizon running cost problem; however, this work does not capture the inverse optimality of the CBF itself.
4.4 Example: Pendulum
We present an example of a pendulum system subjected to disturbance where we demonstrate the robustness of the safety control derived from Proposition 4. We use Proposition 4 to design the following robust safety filter:
Robust min-norm safety filter
$\pi_{S}(x,t)=\arg\min_{u\in U}\left\lVert u-u_{ref}(t)\right\rVert$ (28a)
s.t. $\min_{d\in D}\frac{\partial h}{\partial x}\cdot f(x,u,d)+\gamma h(x)\geq 0,\vspace{-1em}$ (28b)
where $h$ is the chosen CBF. The controller (28) filters a reference control signal $u_{ref}(t)$—when $u_{ref}(t)$ does not satisfy the barrier constraint, it selects a control input $u\in U$ that is closest to $u_{ref}(t)$ that satisfies (28b). Note that from Proposition 4, if we use an FRT value function $V_{\gamma}$ that is differentiable for $h$, the filter is always feasible for $x\in\mathrm{Int}(S)$, and will render the trajectory forward invariant in $\mathrm{Int}(S)$. In the case when the system is affine in control input and disturbance, this safety filter can be implemented as a quadratic program [4, 13].
The dynamics of the pendulum system is given as
$\begin{bmatrix}\dot{x_{1}}\\ \dot{x_{2}}\end{bmatrix}=\begin{bmatrix}x_{2}\\ -\sin{x_{1}}\end{bmatrix}+\begin{bmatrix}0\\ 1\end{bmatrix}u+\begin{bmatrix}0\\ \cos{x_{1}}\end{bmatrix}d,\vspace{-0.5em}$ (29)
where $x_{1}=\theta$, $x_{2}=\dot{\theta}$ are the angle and the angular rate of the pendulum, respectively, $u$ is the applied torque, the control input, and $d$ is the horizontal acceleration applied to the base of the pendulum, the disturbance to the system. $x_{1}=0$ at the released configuration. The desired safety constraint is $X=\{x~{}|~{}0.5\pi\leq x_{1}\leq 2\pi,|x_{2}|\leq 1\}$, constraining both the range of the angle and the angular rate. We set the maximum torque $\bar{u}=\sin{\frac{\pi}{3}}$, so that the maximum torque cannot resist the torque produced by the gravity in the range of $\theta\in[\frac{\pi}{3},\frac{2\pi}{3}]$ and $\theta\in[\frac{4\pi}{3},\frac{5\pi}{3}]$. Then $U$ is set as $[-\bar{u},\bar{u}]$, and $D$ is set as $[-0.1,0.1]$.
Figure 3: (a) The desired safety constraint for the pendulum example is set to be $X=\{x|0.5\pi\leq\theta\leq 2\pi,|\dot{\theta}|\leq 1\}$. The robust control invariant set $S$ is designed by a Bezier fitting to the maximal control invariant set of $X$, computed from the backward reachability analysis. (b) The corresponding target function $h_{S}(X)$ and the computed FRT value function $V_{\gamma}(x)$ with $\gamma=5$.
The reference control is produced by a clipped feedback linearization controller $u_{ref}(t)=\pi_{ref}(\mathrm{x}(t))$, given as
$\pi_{ref}(x)\!=\!\min\!\left\{\max\{\sin{x_{1}}\!-\!k_{1}(x_{1}\!-\!x_{1,\text{t}})\!-\!k_{2}x_{2},-\bar{u}\},\bar{u}\right\},\vspace{-0.5em}$
where the desired target angle $x_{1,\text{t}}$ is set as $-0.2$ for $t\in[0,8]$ and $\pi-0.6$ for $t>8$.
For the design of the target function $h_{S}$, we first verify the maximal control invariant set contained in $X$, using the backward reachability analysis, presented in [17, 38]. Then, a control invariant set with a differentiable boundary, $S$, and its distance function $h_{S}$, is designed by applying Bezier curve fitting [30] to the maximal control invariant set. The resulting $S$ and $h_{S}$ are visualized in Figure 3. We then compute the FRT value function $V_{\gamma}$. The result when $\gamma=5$ is presented in Figure 3, which is differentiable in $\mathrm{Int}(S)$.
[…]
5 Discussion
Although inspired by existing works that establish the connection between reachability and robust control invariance, our work is the first paper that connects forward reachability to the analysis of robust control invariance. Existing works have focused on backward reachability-based formulations that produce the largest robust control invariant set contained in a given desired safety region called the viability kernel. This is done by finding the inevitable (or minimal) backward reachable tube (BRT) [27] of the unsafe region, which becomes the complement of the viability kernel. Various value functions have been proposed to characterize the viability kernels by computing the BRTs, both for finite [13, 26, 28, 19] and infinite time [17, 38, 2, 35] but none of these functions are CBFs, as indicated in Table 1.
The time-varying value functions that characterize finite-horizon BRTs are proposed in [26, 28, 19], where the discount is not needed for the boundedness and continuity of the value function, and the solution uniqueness of the HJ-PDE. The work in [17] has extended these formulations to the infinite-horizon setting (Table 1, first row). However, the value function can be discontinuous and the corresponding HJ-PDE admits non-unique solutions [20]. The work in [2, 38] proposes formulations that resolve these issues for the infinite-horizon setting through the introduction of the discount factor. However, the differential inequality condition on the value function that emerges from the corresponding HJ-PDE differs from the barrier constraint (Table 1, second row). Moreover, the value function flattens to zero in the interior of the control invariant set, thus, there is no non-zero gradient of the value function inside the control invariant set that can be useful for the synthesis of safety control. Thus, this value function is distinct from CBFs, and the authors in [39] extend this formulation to a new notion called Guidance-barrier functions. On the other hand, the work in [13] presents a formulation for the finite-horizon BRT wherein the differential inequality derived from the corresponding HJ-PDE matches the barrier constraint in the CBF definition. However, its extension to infinite horizon might lead to an unbounded and discontinuous value function and non-unique solutions to the corresponding HJ-PDE (Table 1, third row).
Our formulation of forward reachability ensures compliance with the barrier constraint in the CBF definition. Additionally, the value function is both continuous and bounded in $\mathbb{R}^{n}$, while the corresponding HJ-PDE has a unique solution. The central idea behind our approach is the usage of the discount factor backward in time, as $e^{\gamma t}$ where $t<0$, in the definition of the discounted FRT value function (18). In contrast to the discount in BRT formulations leading to the emergence of $-\gamma V_{\gamma}(x)$ in the corresponding HJ-PDEs (Table 1, second row), the usage of discount in this way leads to the emergence of positive $\gamma V_{\gamma}(x)$ term in the FRT-HJ-VI (21), and thus the satisfaction of the barrier constraint (Table 1, last row). Moreover, $e^{\gamma t}$ vanishes as $t$ approaches $\infty$- ∞, thereby ensuring continuous, bounded value functions and the solution uniqueness of the FRT-HJ-VI, resulting from the contraction mapping property outlined in Section 4.2.
An example in Appendix 7.8 illustrates that previous approaches cannot satisfy the barrier constraint, continuity, and boundedness all at the same time, whereas our formulation does.
6 Conclusions
In this study, we have presented a framework that establishes a strong linkage between reachability, control invariance, and Control Barrier Functions (CBFs) through a Hamilton-Jacobi differential game formulation. Two main aspects of our approach are the use of forward reachability concept in lieu of backward reachability, and the incorporation of a discount factor in the value function. These elements induce a contraction in the Bellman backup of the value function, thereby shaping it to satisfy the barrier constraint of the CBFs. Importantly, we note that prior formulations relying on backward reachability were unable to establish this connection between reachability, control invariance, and CBFs. Thus, our work fills a crucial gap in the existing literature, presenting a new perspective on the interplay among these key concepts in the safe control literature through the lens of forward reachability. The emergence of the barrier constraint in the forward reachability formulation opens new avenues for constructing CBFs grounded in reachability analysis.
As we look toward future research avenues, several open questions and challenges emerge. One salient assumption underlying our study is the differentiability of the boundaries of control invariant sets. A deeper understanding of the implications and limitations of this assumption is crucial for broadening the applicability of our results. For example, the relevance of such regularity properties of the boundary of the safe set for stochastic systems is elucidated in [11]. Moreover, the potential of forward reachability, especially in the context of inevitable FRT, has been discussed but not yet fully explored. For instance, our finding of the contraction mapping of the discounted FRT Bellman backup has potential ramifications for learning-based approaches. This property may pave the way for advancements in value-function-based approximate dynamic programming algorithms for safety control [18, 7].
7 Appendix
7.1 Proof of Lemma 7
We define two one-to-one functions: $\rho_{\mathrm{u}}:\mathcal{U}\rightarrow\mathcal{U}^{-}$:
$\rho_{\mathrm{u}}(\mathrm{u})(-t)=\mathrm{u}(t),\quad\forall t\in[0,\infty);\vspace{-0.5em}$ (30)
and $\rho_{\xi_{d}}:\Xi_{d}\rightarrow\Xi_{d}^{-}$:
$\rho_{\xi_{d}}(\xi_{d})[\mathrm{u}^{-}](-t)=\xi_{d}[\rho_{\mathrm{u}}^{-1}(\mathrm{u}^{-})](t),\vspace{-0.5em}$ (31)
for all $\mathrm{u}^{-}\in\mathcal{U}^{-}$ and $t\in[0,\infty)$. Then,
$\mathrm{x}^{-}(-t)=\mathrm{x}(t)\quad\forall t\in[0,\infty),\vspace{-0.5em}$ (32)
where $\mathrm{x}^{-}$ solves (13), and $\mathrm{x}$ solves
$\dot{\mathrm{x}}(t)=-f(\mathrm{x}(t),\rho_{\mathrm{u}}^{-1}(\mathrm{u}^{-})(t),\rho_{\xi_{d}}^{-1}(\Xi_{d}^{-})[\rho_{\mathrm{u}}^{-1}(\mathrm{u}^{-})](t))\vspace{-0.25em}$
for $t>0$, and $\mathrm{x}(0)=x$. Since $\rho_{\mathrm{u}}$ and $\rho_{\xi_{d}}$ are one-to-one, the viability kernel of $\mathrm{Int}(S)^{c}$ under $-f$, defined in Definition 7, is equivalent to the following set:
[…]
$\forall t\in[0,T],\mathrm{x}(t)\in C+B_{\varepsilon}\},\vspace{-0.5em}$ (33)
where $\mathrm{x}$ solves
By [38, Lemma 1], the viability kernel (33) of $\mathrm{Int}(S)^{c}$ is characterized by a particular value function:
[…]
$\mathrm{FRT}(\mathrm{Int}(S))=\{x~{}|~{}V_{\gamma}(x)\neq 0\}.\vspace{-0.5em}$ (35)
Since $h_{S}$ is bounded,
$V_{\gamma}(x)\geq\sup_{\xi_{d}^{-}\in\Xi_{d}^{-}}\inf_{\mathrm{u}^{-}\in\mathcal{U}^{-}}[e^{\gamma t}h_{S}(\mathrm{x}^{-}(t))~{}|~{}t=\-\infty]=0,$ (36)
for all $x\in\mathbb{R}^{n}$. By combining (35) and (36), $\mathrm{FRT}(\mathrm{Int}(S))\!=\!\{x~{}|~{}V_{\gamma}(x)>0\}$.
7.2 Proof of Proposition 2
For $x_{1}\in\mathbb{R}^{n}$ and $\varepsilon>0$, there exists $\hat{\xi}_{d}^{-}\in\Xi_{d}^{-}$ such that
$V_{\gamma}(x_{1})\leq\inf_{\mathrm{u}^{-}\in\mathcal{U}^{-}}J_{\gamma}(x_{1},\mathrm{u}^{-},\hat{\xi}_{d}^{-})+\varepsilon,\vspace{-0.5em}$
where $J_{\gamma}$ is defined in (19). Hence,
$V_{\gamma}(x_{1})\leq J_{\gamma}(x_{1},\mathrm{u}^{-},\hat{\xi}_{d}^{-})+\varepsilon\vspace{-0.5em}$
for any $\mathrm{u}^{-}\in\mathcal{U}^{-}$. For $x_{2}$, there exists $\hat{\mathrm{u}}^{-}\in\mathcal{U}^{-}$ such that
$V_{\gamma}(x_{2})\geq\inf_{\mathrm{u}^{-}\in\mathcal{U}^{-}}J_{\gamma}(x_{2},\mathrm{u}^{-},\hat{\xi}_{d}^{-})\geq J_{\gamma}(x_{2},\hat{\mathrm{u}}^{-},\hat{\xi}_{d}^{-})-\varepsilon.\vspace{-0.5em}$
By combining the above two inequalities, we have
$V_{\gamma}(x_{1})-V_{\gamma}(x_{2})\leq J_{\gamma}(x_{1},\hat{\mathrm{u}}^{-},\hat{\xi}_{d}^{-})-J_{\gamma}(x_{2},\hat{\mathrm{u}}^{-},\hat{\xi}_{d}^{-})+2\varepsilon$
$=\sup_{t\in(-\infty,0]}\!\!e^{\gamma t}h_{S}(\mathrm{x}_{1}^{-}(t))\;-\!\!\sup_{t\in(-\infty,0]}\!\!e^{\gamma t}h_{S}(\mathrm{x}_{2}^{-}(t))+2\varepsilon,\vspace{-0.5em}$
where $\mathrm{x}_{1}^{-}$ solves (13) for $(\hat{\mathrm{u}}^{-},\hat{\xi}_{d}^{-})$ with the terminal state $x_{1}$, and $\mathrm{x}_{2}^{-}$ solves (13) for $(\hat{\mathrm{u}}^{-},\hat{\xi}_{d}^{-})$ with the terminal state $x_{2}$. Since there exists $\hat{t}\in(-\infty,0]$ such that
$\sup_{t\in(-\infty,0]}e^{\gamma t}h_{S}(\mathrm{x}_{1}^{-}(t))\leq e^{\gamma\hat{t}}h_{S}(\mathrm{x}_{1}^{-}(\hat{t}))+\varepsilon,\vspace{-0.5em}$
this implies
$V_{\gamma}(x_{1})-V_{\gamma}(x_{2})\leq e^{\gamma\hat{t}}h_{S}(\mathrm{x}_{1}^{-}(\hat{t}))-e^{\gamma\hat{t}}h_{S}(\mathrm{x}_{2}^{-}(\hat{t}))+3\varepsilon$
$\leq L_{h_{S}}e^{\gamma\hat{t}}e^{-L_{f}\hat{t}}\lVert x_{1}-x_{2}\rVert+3\varepsilon\leq L_{h_{S}}\lVert x_{1}-x_{2}\rVert+3\varepsilon,\vspace{-0.5em}$
where $L_{h_{S}}$ is the Lipschitz constant of $h_{S}$. The second inequality is a result of Gronwall’s inequality, and the third is a result of the condition, $L_{f}<\gamma$. Using a similar argument, we can show $V_{\gamma}(x_{2})-V_{\gamma}(x_{1})\leq L_{h_{S}}\lVert x_{1}-x_{2}\rVert+3\varepsilon$, thus $|V_{\gamma}(x_{1})-V_{\gamma}(x_{2})|\leq L_{h_{S}}\lVert x_{1}-x_{2}\rVert+3\varepsilon$. Since the previous inequality holds for all $\varepsilon>0$, $|V_{\gamma}(x_{1})-V_{\gamma}(x_{2})|\leq L_{h_{S}}\lVert x_{1}-x_{2}\rVert$.
7.3 Proof of Theorem 2
Using the one-to-one mappings $\rho_{\mathrm{u}}$ and $\rho_{\xi_{d}}$ in (30) and (31), the value function $V_{\gamma}$ in (18) can be written as
$V_{\gamma}(x)=\sup_{\xi_{d}\in\Xi_{d}}\inf_{\mathrm{u}\in\mathcal{U}}\sup_{t\in[0,\infty)}e^{-\gamma t}h_{S}(\mathrm{x}(t)),$ (37)
where $\mathrm{x}$ solves the negated dynamics, (34). Then, by adopting the results in [38], we can first present Bellman’s principle of optimality of $V_{\gamma}$:
Theorem 5.
(Dynamic Programming principle [38, Lemma 3]). Suppose $\gamma>0$. For $x\in\mathbb{R}^{n}$,
[…]
$e^{-\gamma T}V_{\gamma}\left(\mathrm{x}^{-}(-T)\right)\!\Big{\}}$
for any $T>0$, where $\mathrm{x}^{-}$ solves (13).
Next, Theorem 2 holds by [38, Lemma 3], since $V_{\gamma}$ rewritten as (37) is the unique viscosity solution to
[…]
$-\min_{u\in U}\max_{d\in D}\frac{\partial V_{\gamma}}{\partial x}\cdot(-f(x,u,d))+\gamma V_{\gamma}(x)\big{\}}$
in $\mathbb{R}^{n}$, which is equivalent to (21).
7.4 Proof of Theorem 3
Define $l(\xi_{d},\mathrm{u},x)\coloneqq\max_{t\in[-T,0]}e^{\gamma t}h_{S}(\mathrm{x}(t))$, and
$l^{i}(\xi_{d},\mathrm{u},x)\coloneqq e^{-\gamma T}V^{i}(\mathrm{x}(-T)),\vspace{-0.5em}$
for $i=1,2$. Then,
$B_{T}[V^{i}]=\sup_{\xi_{d}\in\Xi_{d}}\inf_{\mathrm{u}\in\mathcal{U}}\max\{l(\xi_{d},\mathrm{u}),l^{i}(\xi_{d},\mathrm{u})\}.\vspace{-0.5em}$
[…]
$B_{T}[V^{1}](x)-B_{T}[V^{2}](x)$
$\;<2\varepsilon+\max\{l(\bar{\xi}_{d},\bar{\mathrm{u}}),l^{1}(\bar{\xi}_{d},\bar{\mathrm{u}})\}-\max\{l(\bar{\xi}_{d},\bar{\mathrm{u}}),l^{2}(\bar{\xi}_{d},\bar{\mathrm{u}})\}$
$\;\leq 2\varepsilon+\lvert l^{1}(\bar{\xi}_{d},\bar{\mathrm{u}})-l^{2}(\bar{\xi}_{d},\bar{\mathrm{u}})\rvert$
$\;\leq 2\varepsilon+e^{-\gamma T}\max_{x\in\mathbb{R}^{n}}\lvert V^{1}(x)-V^{2}(x)\rvert\vspace{-1em}$
The second inequality holds since, for all $a,b,c\in\mathbb{R}$, $|\max\{a,b\}-\max\{a,c\}|\leq|b-c|$. Since the above inequality holds for all $x\in\mathbb{R}^{n}$ and $\varepsilon>0$,
$\lVert B_{T}[V^{1}]-B_{T}[V^{2}]\rVert_{L^{\infty}(\mathbb{R}^{n})}\leq e^{-\gamma T}\lVert V^{1}-V^{2}\rVert_{L^{\infty}(\mathbb{R}^{n})}\vspace{-0.5em}$
Since $V_{\gamma}$ is a fixed-point solution for all $T>0$, the Banach’s contraction mapping theorem [15, Chapter 9.2] implies that $V_{\gamma}$ is the unique fixed-point solution to $B_{T}[V_{\gamma}](x)=V_{\gamma}(x)$ for all $T>0$. In addition, we have
$\lVert B_{T}[V]-V_{\gamma}\rVert_{L^{\infty}(\mathbb{R}^{n})}\leq e^{-\gamma T}\lVert V-V_{\gamma}\rVert_{L^{\infty}(\mathbb{R}^{n})}\vspace{-0.5em}$
for all $V\!\in\!\textnormal{BUC}(\mathbb{R}^{n})$, thus, $\lim_{T\rightarrow\infty}B_{T}[V]=V_{\gamma}$.
7.5 Computation methods for $V_{\gamma}$
First, the following lemma presents a finite-horizon HJ equation for the computation of $V_{\gamma}$ using the Bellman operator $B_{T}$.
Lemma 8 (Finite horizon HJ-PDE for the computation of $V_{\gamma}$).
For a given initial value function candidate $V^{0}\in\textnormal{BUC}(\mathbb{R}^{n})$, let $W:[0,T]\times\mathbb{R}^{n}\rightarrow\mathbb{R}$ be the unique viscosity solution to the following initial-value HJ-PDE
$W(0,x)=\max\{h_{S}(x),V^{0}(x)\},\text{ for }x\in\mathbb{R}^{n},$ (39)
$0=\min\bigg{\{}W(t,x)-h_{S}(x),$ (40)
$\quad\quad\frac{\partial W}{\partial t}+\max_{u}\min_{d}\frac{\partial W}{\partial x}\!\cdot\!f(x,u,d)+\gamma W(t,x)\bigg{\}}$
for $(t,x)\in(0,T)\times\mathbb{R}^{n}$. Then, $W(T,x)\equiv B_{T}[V^{0}](x)$.
Proof.
We will derive the HJ equation for another value function $W^{+}$ defined below, and then replace $W^{+}$ by $W$. Define $W^{+}:[-T,0]\times\mathbb{R}^{n}\rightarrow\mathbb{R}$
[…]
$e^{\gamma t}(-V_{0}(\mathrm{x}(0)))\bigg{\}},$ (41)
where $\mathrm{x}$ solves (34). Then, $W(T,x)=-W^{+}(-T,x)$, and
[…]
We adopt the results in [13] for the rest of the proof. The value function in [13] is
$\inf_{\xi_{d}\in\xi_{d}}\sup_{\mathrm{u}}\min_{s\in[t,0]}e^{-\gamma(s-t)}(-h_{S}(\mathrm{x}(s))),\vspace{-0.5em}$
which is the exactly same as (41) except the second term of the minimization operation: $e^{\gamma t}(-V_{0}((\mathrm{x}(0)))$. This term affects the terminal condition of $W^{+}$ but not the dynamic programming principle. Thus, $W^{+}$ and the value function in [13] solves the same dynamic programming principle, but their terminal conditions are different. Note that we assume $\gamma>0$, but [13] assumes $\gamma\leq 0$. However, the sign of $\gamma$ does not affect any arguments in [13]’s lemmas, theorems.
Using similar arguments as in the proof of [13, Theorem 2], we get
[…]
$e^{-\gamma\delta}W^{+}(t+\delta,\mathrm{x}(t+\delta))\bigg{\}}.\vspace{-0.5em}$
Then, [13, Theorem 3] implies that $W^{+}$ is the unique viscosity solution to the terminal-value HJ equation:
$W^{+}(0,x)=-\max\{h_{S}(x),V_{0}(x)\}\quad\text{on }\{t=0\}\times\mathbb{R}^{n},$
$0=\min\Big{\{}-\!h_{S}(x)\!-\!W^{+}(t,x),$
$\quad\quad\quad\frac{\partial W^{+}}{\partial t}+\max_{u}\min_{d}\frac{\partial W^{+}}{\partial x}\cdot(-f)(x,u,d)-\gamma W^{+}(t,x)\Big{\}},\vspace{-1em}$
in $(-T,0)\times\mathbb{R}^{n}$. By applying (42), we get the conclusion that $W$ is the unique viscosity solution to (40). ∎
In Lemma 8, any $V^{0}\in\textnormal{BUC}(\mathbb{R}^{n})$ works for the computation of $V_{\gamma}$; for instance, a straightforward choice of $V^{0}$ can be $h_{S}$. As $T\rightarrow\infty$, $\frac{\partial W}{\partial t}$ vanishes to 0 for all $x\in\mathbb{R}^{n}$.
Combining Theorem 3 and Lemma 8, we have
$\lim_{T\rightarrow\infty}B_{T}[V^{0}]=\lim_{T\rightarrow\infty}W(T,x)=V_{\gamma}(x).$ (43)
The PDE (40) can be numerically solved forward in time from the initial condition (39), by using well-established time-dependent level-set methods [29].
Theorem 3 also enables other numerical schemes that are based on time discretization, like value iteration, to produce an accurate solution of $V_{\gamma}$. The following corollary of Theorem 3 provides the guarantee that the value iteration with any initial guess of $V^{0}\in\textnormal{BUC}(\mathbb{R}^{n})$ will converge to $V_{\gamma}$ with a Q-linear convergence rate specified by (44). For a given time step size $\Delta t$, the semi-Lagrangian approximation can be applied to the exact Bellman backup operator in (22) for its numerical approximation, and the resulting value function will converge to $V_{\gamma}$ when $\Delta t\rightarrow 0$ [2].
Corollary 3 (Value Iteration).
For any $V^{0}\in\text{BUC}(\mathbb{R}^{n})$ and a time step $\Delta t>0$, define the sequence $\{V^{k}\}_{k=0}^{\infty}$ by an iteration $V^{k}:=B_{\Delta t}[V^{k-1}]$ for $k\in\mathbb{N}$. Then,
$\frac{\lVert V^{k+1}-V_{\gamma}\rVert_{\infty}}{\lVert V^{k}-V_{\gamma}\rVert_{\infty}}\leq e^{-\gamma\Delta t}<1,$ (44)
and thus, $\lim_{k\rightarrow\infty}V^{k}=V_{\gamma}$.
Proof.
This is a direct outcome of Theorem 3. ∎
7.6 Proof of Proposition 4
(i) At $x\in\mathrm{Int}(S)$ where $V_{\gamma}$ is differentiable, the FRT-HJ-VI (21) implies that $K_{\gamma}$ is non-empty.
(ii) For any control policy $\pi=\pi(x)\in K_{\gamma}(x)$, where $V_{\gamma}$ is differentiable, consider the following equation for $V^{\pi}_{\gamma}$:
$0\!=\!\min\Big{\{}\!\ V^{\pi}_{\gamma}(x)-h_{S}(x),\min_{d}\frac{\partial V_{\gamma}}{\partial x}\cdot f(x,\pi(x),d)+\gamma V^{\pi}_{\gamma}(x)\Big{\}}.$
For each $x\in\mathbb{R}^{n}$, $\min\{y-h_{S}(x),\min_{d}\frac{\partial V_{\gamma}}{\partial x}\cdot f(x,\pi(x),d)+\gamma y\}$ is monotonically increasing in $y$, so the equation (45) has a unique solution. Also, from the FRT-HJ-VI (21),
The last inequality holds since $V_{\gamma}-h_{S}\geq 0$ from (21) and $\min_{d}\frac{\partial V_{\gamma}}{\partial x}\cdot f(x,\pi(x),d)+\gamma V_{\gamma}(x)\geq 0$ since $\pi(x)\in K_{\gamma}(x)$. Equation (46) and the uniqueness of (45) imply $V_{\gamma}\equiv V^{\pi}_{\gamma}$ for any $\pi$. By replacing $V_{\gamma}$ by $V^{\pi}_{\gamma}$ in (45),
$0\!=\!\min\Big{\{}\!\ V^{\pi}_{\gamma}(x)\!-\!h_{S}(x),\min_{d}\frac{\partial V^{\pi}_{\gamma}}{\partial x}\cdot f(x,\pi(x),d)+\gamma V^{\pi}_{\gamma}(x)\Big{\}}$.
The solution to the above PDE can be considered as the value function (19) under $\pi(x)$ and worst-case disturbance, and since $V_{\gamma}\equiv V^{\pi}_{\gamma}$, we conclude that any control $u\in K_{\gamma}(x)$ is an optimal control for the zero-sum game value $V_{\gamma}$ in (18).
7.7 Proof of Theorem 4
We will show the two statements as follows.
  1. 1.
    $\forall v\in C^{\infty}(\mathbb{R}^{n})$ such that $V_{\gamma}-v$ has a local minimum at $x_{0}\in\mathbb{R}^{n}$ and $V_{\gamma}(x_{0})=v(x_{0})$,
    $0\leq\min\Big{\{}v(x_{0})-h(x_{0}),\max_{u\in U}\min_{d\in D}\frac{\partial v}{\partial x}(x_{0})\cdot f(x_{0},u,d)+\gamma v(x_{0})\Big{\}}.$
  2. 2.
    $\forall v\in C^{\infty}(\mathbb{R}^{n})$ such that $V_{\gamma}-v$ has a local maximum at $x_{0}\in\mathbb{R}^{n}$ and $V_{\gamma}(x_{0})=v(x_{0})$,
    $0\geq\min\Big{\{}v(x_{0})-h(x_{0}),\max_{u\in U}\min_{d\in D}\frac{\partial v}{\partial x}(x_{0})\cdot f(x_{0},u,d)+\gamma v(x_{0})\Big{\}}.$
Case 1. $V_{\gamma}(x_{0})\!=\!h(x_{0})\!>\!0$: By the continuity of $h$, there exists $\varepsilon>0$ such that $V_{\gamma}(y)\!=\!h(y)$ for all $y\!\in\!B_{\varepsilon}(x_{0})$. Thus, the gradient of $V_{\gamma}$ at $x_{0}$ exists: $\frac{\partial V_{\gamma}}{\partial x}(x_{0})=\frac{\partial h}{\partial x}(x_{0})$, so for any $v$ such that $V_{\gamma}-v$ has either a local minimum or a local maximum at $x_{0}$, $\frac{\partial v}{\partial x}(x_{0})=\frac{\partial h}{\partial x}(x_{0})$. From (26),
$\max_{u\in U}\min_{d\in D}\frac{\partial v}{\partial x}(x_{0})\cdot f(x_{0},u,d)+\gamma v(x_{0})=0.\vspace{-0.5em}$ (47)
Therefore, Statements 1) and 2) hold in this case.
Case 2. $V_{\gamma}(x_{0})\!=\!0\!>\!h(x_{0})$: By the continuity of $h$, there exists $\varepsilon>0$ such that $V_{\gamma}(y)=0$ for all $y\in B_{\varepsilon}(x_{0})$. This implies that the gradient of $V_{\gamma}$ at $x_{0}$ is $0\in\mathbb{R}^{n}$, so for any $v$ such that $V_{\gamma}-v$ has either a local minimum or a local maximum at $x_{0}$, $\frac{\partial v}{\partial x}(x_{0})=0$. Thus, (47) holds. Therefore, Statements 1) and 2) hold in this case.
Case 3. $V_{\gamma}(x_{0})\!=\!0\!=\!h(x_{0})$: From $v(x_{0})-h(x_{0})\!=\!0$, it is trivial that 2) holds, and we focus on the proof of 1). Since $V_{\gamma}-v$ has a local minimum at $x_{0}$, $\frac{\partial v}{\partial x}(x_{0})\in\partial^{-}V_{\gamma}(x_{0})$, where $\partial^{-}V_{\gamma}(x_{0})$ is the sub-differential, which is determined as $\partial^{-}V_{\gamma}(x_{0})=\text{conv}\left(\{0\}\cup\Big{\{}\frac{\partial h}{\partial x}(x_{0})\Big{\}}\right)$, where conv is a convex-hull operator. Thus, $\frac{\partial v}{\partial x}(x_{0})=\alpha\frac{\partial h}{\partial x}(x_{0})$ for some $\alpha\in[0,1]$. Thus, from (26) and $v(x_{0})=0$, (47) holds and therefore, 1) holds.
7.8 1D example for comparison of methods in Table 1
We consider a simple one-dimensional system:
$\dot{\mathrm{x}}(t)=\mathrm{x}(t)+\mathrm{u}(t),\;\mathrm{x}(0)=x,\vspace{-0.5em}$ (48)
with $u\in U=[-1,1]$, and the state domain $x\in[0,\infty)$. We do not introduce disturbance for simplicity. We consider the set $S=[0,2]$, and choose $h_{S}$ as $h_{S}(x)=\max\{2\!-\!x,-2\}$. Basically, it is a distance function cut off at the absolute value of $2$ (Figure 5 grey). We compare the results of the three backward reachability formulations studied in the previous literature [17, 2, 38, 13, 35] with our forward reachability formulation, as summarized in Table 1. We use $\gamma=2$ in all formulations. Note that the chosen $S$ is not control invariant, thus, the example is chosen not for the reachability analysis of control invariant sets, but to study the boundedness, continuity, and the solution uniqueness of the resulting value functions. The maximal control invariant set contained in $S$, is $[0,1]$, since at $x=1$, $\dot{x}$ can be maintained 0 by selecting the saturated control input $u=-1$, but for every $x$ exceeding 1, $\dot{x}>0$ for any admissible $u\in U$.
Figure 5: Various reachability value functions, summarized in Table 1, for the 1D example in (48). Only the formulations with the discounted factor (magenta and blue), including the one proposed in this work produce value functions that are bounded and continuous.
Backward Reachable Tube (BRT) without discount factor [17, 20]:
$V(x):=\underset{{\xi_{d}\in\Xi_{d}}}{\inf}\;\underset{\mathrm{u}\in\mathcal{U}}{\sup}\;\underset{t\in[0,\infty)}{\inf}h_{S}(\mathrm{x}(t)),\vspace{-0.5em}$ (49)
which characterizes the viability kernel of $S$ as $\{x|V(x)\geq 0\}$, which can be seen in Figure 5 (purple). The value function is discontinuous at $x=1$. Moreover, the corresponding HJ-PDE, given as
$0=\min\Big{\{}\ h_{S}(x)-V(x),\max_{u}\min_{d}\frac{\partial V}{\partial x}\cdot f(x,u,d)\Big{\}}$
admits non-unique solutions, for instance, $V(x)\equiv-2$ in this example is also a valid viscosity solution to (50).
BRT with discount factor [2, 38]:
$V(x):=\underset{{\xi_{d}\in\Xi_{d}}}{\inf}\;\underset{\mathrm{u}\in\mathcal{U}}{\sup}\;\underset{t\in[0,\infty)}{\inf}e^{-\gamma t}h_{S}(\mathrm{x}(t)).\vspace{-0.5em}$ (51)
In this case, as can be seen in Figure 5 (green), the value function is continuous and bounded. However, the value function is flat inside the viability kernel, which is characterized as $\{x|V(x)=0\}$.
Infinite horizon CBVF [13, 35]:
$V(x):=\underset{{\xi_{d}\in\Xi_{d}}}{\inf}\;\underset{\mathrm{u}\in\mathcal{U}}{\sup}\;\underset{t\in[0,\infty)}{\inf}e^{\gamma t}h_{S}(\mathrm{x}(t)).\vspace{-0.5em}$ (52)
Notice the flip of the sign in the factor of the exponential term, compared to (51). This formulation results in an HJ-PDE whose differential inequality matches the form of the barrier constraint. However, it results in discontinuity and unboundedness of the value function, as can be seen in Figure 5 (orange).
Our formulation: The value function is defined in (18), which is bounded and continuous, as can be seen in Figure 5 (blue). Also, this formulation admits a unique solution to the corresponding HJ-PDE in (21), and the differential inequality in the PDE matches the form of the barrier constraint. Note that in this example, $\mathrm{FRT}(\mathrm{Int}(S))$ is $[0,\infty)$, thus, in Figure 5, $V(x)>0$ everywhere.
Acknowledgments
We thank Professors Helena Frankowska and Lawrence C. Evans for insightful discussions.